17. If one randomly picks two different months from the Tornado data list. What is the probability that they both are Summer months (June, July, August).

  • a. 1/12
  • b. 2/12
  • c. 3/66

18. If one randomly picks two different months from the Tornado data list. What is the probability that their Correlation will be above 0.5. Hint: Check the correlation table.

  • a. 2/66
  • b. 2/33
  • c. 1/18

19. If a person rolls two dice, what is the probability that the sum of the numbers is equal to 7?

  • a. 1/18
  • b. 1/6
  • c. 1/36

20. If a person tosses three coins what is the probability that he will not have any Heads?

  • a. 1/8
  • b. 2/8
  • c. 1/2
Answer #1

17) total possible outcomes = 12c2 = 66

Pairs of summer months possible = 3

Required probability = 3/66

Option (c)

19) possible ways to get the sum of 7 = (1,6) (6,1) (2,5) (5,2) (3,4) (4,3) = 6 possible outcomes

Total possible outcomes when a pair of dice rolled = 36

Required probability = 6/36) = 1/6

Option (b)

20) possible outcomes = HHH, TTT, TTH, HTT, HHT, THH, THT, HTH. = 8 possible outcomes

Their's only 1 outcome which don't has any H and that is TTT = 1 outcome

Required probability = 1/8

Option (a)